two groups of tourists traveled by automobiles. one group averaged 45 kilometers per hour and the other 35 kilometers per hour. the faster group traveled 30 kilometers farther than the slower group and made the trip in two hours less time. how long did each party travel?
Answer by lwsshak3(11628)

two groups of tourists traveled by automobiles. one group averaged 45 kilometers per hour and the other 35 kilometers per hour. the faster group traveled 30 kilometers farther than the slower group and made the trip in two hours less time. how long did each party travel?
**
let x=distance traveled by slower group
x+30=distance traveled by faster group
time of travel=distance/speed
..
(x/35)-(x+30)/45=2
LCD: 35*45
(45x)-35(x+30)=2*35*45
45x-35x-1050=3150
10x=4200
x=420 km
ans:
slower group traveled for 420/35=12 hrs
slower group traveled for 420+30/45=10 hrs